# Transport alpha divergences

### v2

@@ 4. Hessian structures of entropy in Wasserstein-$2$ space
In this section, we formulate the Hessian structures in Wasserstein-$2$ space on one-dimensional sample space. In particular, we derive the 3-symmetric tensor from the third order derivatives of Boltzmann–Shannon entropy in Wasserstein-$2$ space.
4.1. Review of Hessian metric in Wasserstein-$2$ space
@@
For any constant $c\in\mathbb{R}$, if $\Phi\in T_{p}^{*}\mathcal{P}_{o}(\Omega)$, then $\Phi(x)+c\in T_{p}^{*}\mathcal{P}_{o}(\Omega)$.
Define an inner product $g_{\mathrm{W}}\colon\mathcal{P}_{o}(\Omega)\times T_{p}\mathcal{P}_{o}(\Omega)\times T_{p}\mathcal{P}_{o}(\Omega)\rightarrow\mathbb{R}$ by
@@
where $\sigma_{i}(x)=-\partial_{x}\big{(}p(x)\Phi^{\prime}_{i}(x)\big{)}$, with $\sigma_{i}\in T_{p}\mathcal{P}_{o}(\Omega)$ and $\Phi_{i}\in T^{*}_{p}\mathcal{P}_{o}(\Omega)$, for $i=1,2$. Thus, $(\mathcal{P}(\Omega),g_{\mathrm{W}})$ forms an infinite-dimensional Riemannian manifold in probability density space. In literature, it is often called density manifold [14] or Wasserstein-$2$ space [28].
@@
$g_{\mathrm{H}}(p)(\sigma_{1},\sigma_{2}):=-\textrm{Hess}_{\mathrm{W}}\mathcal{H}(\sigma_{1},\sigma_{2}):=\int_{\Omega}\Phi^{\prime\prime}_{1}(x)\cdot\Phi^{\prime\prime}_{2}(x)p(x)dx,$ (15)
@@ 4.2. Transport 3–symmetric tensor
We are now ready to formulate the third derivative of entropy $\mathcal{H}(p)$ in Wasserstein-$2$ space. It is a three form, or 3–symmetric tensor in $(\mathcal{P}(\Omega),g_{\mathrm{W}})$.
@@
We finish this section by representing the Taylor expansions of transport alpha divergences, using Hessian structures $(\mathcal{P}_{o}(\Omega),g_{\mathrm{H}},T_{\mathrm{H}})$.
@@
$\begin{split}\mathrm{D}_{\mathrm{T},\alpha}(p\|q)=&\frac{1}{2}g_{\mathrm{H}}(q)(\sigma,\sigma)+\frac{\alpha-3}{6}T_{\mathrm{H}}(q)(\sigma,\sigma,\sigma)+{O(\int_{\Omega}|\Phi^{\prime\prime}(x)|^{4}q(x)dx)}.\end{split}$
@@
$\int_{\Omega}(\Phi^{\prime\prime}(x))^{k}p(x)dx=\int_{0}^{1}(\frac{Q^{\prime}_{p}(u)}{Q^{\prime}_{q}(u)}-1)^{k}du.$
@@
We note that $\Gamma_{1}$, $\Gamma_{2}$ are often called Gamma one and Gamma two operators, which are firstly introduced by Bakry–Émery [8] to study the Ricci curvature lower bound on a sample space. For simplicity of presentation, we only show them in one-dimensional sample space. The iterative Gamma two calculus connects with second-order derivatives of entropy in Wasserstein-$2$ space [8, 30] with generalizations [15]. Here, we present a “third-order” Gamma calculus to formulate the third derivatives in Wasserstein-$2$ space, namely transport 3-symmetric tensor. We will study geometric calculations of transport-3 symmetric tensors in high-dimensional spaces in future works. Following [15, 16], we expect that information geometry and Gamma three operators are tools in studying generalized divergences in Wasserstein-$2$ type spaces.
@@ 5. Examples
In machine learning applications [4], a generative model is defined as follows. Consider a latent random variable $Z\sim p_{\mathrm{ref}}$, where $p_{\mathrm{ref}}\in\mathcal{P}(\Omega)$ is a given reference measure. Denote an invertible map function $G\colon\Omega\times\Theta\rightarrow\Omega$, where $\Theta\subset\mathbb{R}^{n}$ is a parameter space. Then
@@
Proposition 5 (Transport alpha divergence in generative models).
@@
where we need to assume that $\mathbb{E}_{Z\sim p_{\mathrm{ref}}}|G(Z,\theta)|^{2}<+\infty$, for $\theta=\theta_{X}$ or $\theta_{Y}$.
@@
Suppose $G$ is a linear mapping function such that
@@
We remark that the Cauchy distribution does not have the finite second moment. And the Wasserstein-$2$ distance is not well defined, such that $W_{2}(p(\cdot,\gamma_{1}),p(\cdot,\gamma_{2}))=+\infty$.
@@ 6. Discussion
This paper proposes a class of transport alpha divergences, one-parameter variation of transport KL divergence and transport Hessian distance. They are connected with Hessian metrics and 33–symmetric tensors of the negative Boltzmann-Shannon entropy in Wasserstein-$2$ space. We provide several analytical examples in one-dimensional probability densities, including generative models.
It is worth mentioning that the quantile density functions (QDFs) have been applied in statistical learning problems [29]. The quantile density functions measure densities’ shape up to any constant shifting. The transport alpha divergence provides a class of functionals for measuring the difference from QDFs, i.e., Jacobi functions of mapping functions. In future work, we shall study transport alpha divergences in high dimensional probability densities [10, 19]. This direction includes analysis, dualities, invariance properties, and optimization algorithms of transport mapping-related divergence functionals. In particular, systemic geometric calculations for Hessian structures in Wasserstein-$2$ spaces $(\mathcal{P}_{o}(\Omega),g_{\mathrm{H}},T_{\mathrm{H}})$ will be investigated; see related studies in [9, 19, 20, 26]. The systemic understanding of three symmetric tensors and alpha connections rely on both studies in optimal transport and information geometry, which is not a simple direction; see related studies in [5, 6, 24]. We expect that the convexity analysis and approximations in transport Hessian structures serve the mathematical foundations of artificial intelligence, in particular for studying new optimization problems or approximations in generative models.
Acknowledgements. W. Li’s work is supported by AFOSR YIP award No. FA9550-23-1-0087, NSF RTG: 2038080, and NSF DMS: 2245097.
@@
In this section, we present some necessary proofs for this paper.
@@ 6.1. Derivatives in Wasserstein-$2$ space
We recall that the geodesics in $(\mathcal{P}_{o}(\Omega),g_{\mathrm{W}})$ satisfies
$\left\{\begin{split}&\partial_{t}p(t,x)+\partial_{x}(p(t,x)\partial_{x}\Phi(t,x))=0\\ &\partial_{t}\Phi(t,x)+\frac{1}{2}|\partial_{x}\Phi(t,x)|^{2}=0,\end{split}\right$.
where $p(0,x)=p(x)$ and $\partial_{t}p(0,x)=\sigma(x)=-\partial_{x}(p(x)\Phi^{\prime}(x))$. We prove the result by induction. When $n=1$, we have
@@
These geometric formulas are derived based on the Riemannian Levi-Civita connection in density manifold $(\mathcal{P}_{o}(\Omega),g_{\mathrm{W}})$. They formulate classical Gamma calculuses; see details in [15, 16, 30]. We leave studies of high-order derivatives of entropy in $(\mathcal{P}_{o}(\Omega),g_{\mathrm{W}})$ with high dimensional sample spaces in future works.
@@ 6.2. Transport Hessian distance
Here the infimum is taken among all smooth density paths $p\colon[0,1]\times\Omega\rightarrow\mathbb{R}$, which connects both initial and terminal time probability density functions $q$, $p\in\mathcal{P}(\Omega)$. Interestingly, there is a closed-form update for the variational problem (16).
@@
$\mathrm{Dist}_{\mathrm{H}}(p,q)^{2}=\int_{0}^{1}|\log\frac{Q_{p}^{\prime}(u)}{Q_{q}^{\prime}(u)}|^{2}du=2\mathrm{D}_{\mathrm{T},0}(p\|q).$
@@
where the infimum is taken among all smooth mapping functions $T\colon[0,1]\times\Omega\rightarrow\Omega$ with $T(0,x)=x$ and $T(1,x)=T(x)$. Thus,
@@
where the infimum is taken among all paths $Q^{\prime}_{p_{t}}\colon[0,1]^{2}\rightarrow\mathbb{R}$ with fixed initial and terminal time conditions. By using the Euler-Lagrange equation for variable $Q^{\prime}_{p_{t}}$, we show that the geodesic equation in transport Hessian metric satisfies
@@ 6.3. Transport KL divergence
We also briefly present the motivation of transport KL divergence. Consider a functional $\mathcal{F}\colon\mathcal{P}\rightarrow\mathbb{R}$. We define a class of Bregman divergences in Wasserstein-2 space
@@
where $p$, $q\in\mathcal{P}(\Omega)$, $\frac{\delta}{\delta q(x)}$ is the $L^{2}$ first variation w.r.t. $q(x)$, and $T_{\#}q=p$. We refer $\mathrm{D}_{\mathrm{T},\mathcal{F}}$ as the transport Bregman divergence. If $\mathcal{F}$ is a second moment functional, i.e. $\mathcal{F}(p)=\int_{\Omega}|x|^{2}p(x)dx$, then $\mathrm{D}_{\mathrm{T},\mathcal{F}}$ forms the Wasserstein-2 distance. If $\mathcal{F}(p)=-\mathcal{H}(p)=\int_{\Omega}p(x)\log p(x)dx$, then
@@
where we use the fact that $q(x)\partial_{x}\log q(x)=\partial_{x}q(x)$ in the third equality, and apply the integration by parts in the last equality. From now on, we name $\mathrm{D}_{\mathrm{TKL}}$ the transport KL divergence. Again, by the change of variable (18), we note that $\mathrm{D}_{\mathrm{TKL}}$ can be formulate in terms of quantile density functions:
@@
From this formulation, we observe that $\mathrm{D}_{\mathrm{TKL}}$ is an Itakura–Saito type divergence in term of mapping functions or quantile density functions.
@@ 6.4. Validation of canonical divergences
We follow the idea of proof in [7], with using the change of variable (18) and transport Hessian metric $g_{\mathrm{H}}$. We denote a monotone function $\hat{T}(x)=T^{-1}(x)$, such that $(T^{-1})_{\#}p=q$. Denote $\hat{T}\colon[0,1]\times\Omega\rightarrow\Omega$ satisfying Proposition 4 with $\alpha=-1$, such that
@@
$\begin{split}g_{\mathrm{H}}(\partial_{t}r,\partial_{t}r)=&\int_{\Omega}|\partial_{t}\partial_{x}\hat{T}(t,x)\frac{1}{\partial_{x}\hat{T}(t,x)}|^{2}{q(x)}dx.\end{split}$
@@
$\begin{split}-\frac{d^{2}}{dt^{2}}\mathcal{H}(r(t,\cdot))=&\int_{\Omega}\frac{d^{2}}{dt^{2}}\log\frac{p(x)}{\partial_{x}\hat{T}(t,x)}p(x)dx\\ =&\int_{\Omega}\Big{(}|\partial_{t}\partial_{x}\hat{T}(t,x)\frac{1}{\partial_{x}\hat{T}(t,x)}|^{2}-\frac{\partial_{tt}\partial_{x}\hat{T}(t,x)}{\partial_{x}\hat{T}(t,x)}\Big{)}{q(x)}dx\\ =&g_{\mathrm{H}}(\partial_{t}r,\partial_{t}r).\end{split}$

### v3

@@ 4. Hessian structures of entropy in Wasserstein-$2$ space
In this section, we formulate the Hessian structures in Wasserstein-$2$ space on one-dimensional sample space. In particular, we derive the 3-symmetric tensor from the third order derivatives of negative Boltzmann–Shannon entropy in Wasserstein-$2$ space.
4.1. Review
@@
For any constant $c\in\mathbb{R}$, if $\Phi\in T_{p}^{*}\mathcal{P}_{o}(\Omega)$, then $\Phi(x)+c\in T_{p}^{*}\mathcal{P}_{o}(\Omega)$. Define an inner product $g_{\mathrm{W}}\colon\mathcal{P}_{o}(\Omega)\times T_{p}\mathcal{P}_{o}(\Omega)\times T_{p}\mathcal{P}_{o}(\Omega)\rightarrow\mathbb{R}$ by
@@
where $\sigma_{i}(x)=-\partial_{x}\big(p(x)\Phi^{\prime}_{i}(x)\big)$, with $\sigma_{i}\in T_{p}\mathcal{P}_{o}(\Omega)$ and $\Phi_{i}\in T^{*}_{p}\mathcal{P}_{o}(\Omega)$, for $i=1,2$. Thus, $(\mathcal{P}(\Omega),g_{\mathrm{W}})$ satisfies an infinite-dimensional Riemannian manifold in probability density space. In literature, $(\mathcal{P}(\Omega),g_{\mathrm{W}})$ is often called density manifold [15] or Wasserstein-$2$ space [28].
@@
$g_{\mathrm{H}}(p)(\sigma_{1},\sigma_{2}):=-\textrm{Hess}_{\mathrm{W}}\mathcal{H}(p)(\sigma_{1},\sigma_{2}):=\int_{\Omega}\Phi^{\prime\prime}_{1}(x)\cdot\Phi^{\prime\prime}_{2}(x)p(x)dx,$ (15)
@@ 4.2. Transport 3–symmetric tensor
We are now ready to formulate the third derivative of negative entropy $\mathcal{H}(p)$ in Wasserstein-$2$ space. It is a three form, or 3–symmetric tensor in $(\mathcal{P}(\Omega),g_{\mathrm{W}})$.
@@
We finish this section by representing Taylor expansions of transport alpha divergences, using the Hessian structure $(\mathcal{P}_{o}(\Omega),g_{\mathrm{H}},T_{\mathrm{H}})$.
@@
$\begin{split}\mathrm{D}_{\mathrm{T},\alpha}(p\|q)=&\frac{1}{2}g_{\mathrm{H}}(q)(\sigma,\sigma)+\frac{\alpha-3}{6}T_{\mathrm{H}}(q)(\sigma,\sigma,\sigma)+{O\left(\int_{\Omega}|\Phi^{\prime\prime}(x)|^{4}q(x)dx\right)}.\end{split}$
@@
$\int_{\Omega}(\Phi^{\prime\prime}(x))^{k}p(x)dx=\int_{0}^{1}\left(\frac{Q^{\prime}_{p}(u)}{Q^{\prime}_{q}(u)}-1\right)^{k}du.$
@@
We note that $\Gamma_{1}$, $\Gamma_{2}$ are often called Gamma one and Gamma two operators, which are firstly introduced by Bakry–Émery [8] to study the Ricci curvature lower bound on a sample space. Here we only show them in one-dimensional sample space. The iterative Gamma two calculus connects with second-order derivatives of entropy in Wasserstein-$2$ space [8, 30] with generalizations [16]. Here, we present a “third-order” Gamma calculus to formulate the third derivatives in Wasserstein-$2$ space, namely transport 3-symmetric tensor. We will study geometric calculations of transport-3 symmetric tensors in high-dimensional spaces in future works. Following [16, 17], we expect that the information geometry method and Gamma three operators are tools in studying generalized divergences in high dimensional Wasserstein-$2$ spaces.
@@ 5. Examples
In machine learning applications [4], a generative model is defined as follows. Consider a latent random variable $Z\sim p_{\mathrm{ref}}$, where $p_{\mathrm{ref}}\in\mathcal{P}(\Omega)$ is a given reference measure. Denote a map function $G\colon\Omega\times\Theta\rightarrow\Omega$, where $\Theta\subset\mathbb{R}^{n}$ is a parameter space. We also assume that $G(\cdot,\theta)$ is a monotone mapping for all $\theta\in\Theta$. Then
@@
Proposition 5 (Transport alpha divergence in one dimensional distributions).
@@
where we need to assume that $\mathbb{E}_{Z\sim p_{\mathrm{ref}}}|G(Z,\theta)|^{2}<+\infty$, for $\theta=\theta_{X}$ or $\theta_{Y}$. We note that if $\alpha\neq 0$, the transport alpha divergence is not symmetric on parameters $\theta_{X}$, $\theta_{Y}$. This is in contrast with the Wasserstein-2 distance.
Remark 7.
One may apply the transport alpha divergence to measure the closeness between two neural network parameters. The other application is to apply transport alpha divergences as loss functions for inference problems. Some related studies have been conducted for the quantile density function in statistics [29]. However, in general high-dimensional sample spaces, the construction of transport alpha divergences, with the selection of parameter alpha in inference problems, are unclear, which are left for the future work.
Remark 8.
The proposed transport alpha divergence is not intended as a replacement formula for the Wasserstein-$2$ distance. Instead, it modifies the geometric structure of the density manifold. While the local expansion of the Wasserstein-$2$ distance yields only the Riemannian metric (the Wasserstein information matrix), the Taylor expansion of the transport alpha divergence contains an additional third-order operator, which beyonds the transport Hessian metric. At present, the optimization properties of the divergence as a loss function are not fully understood. The systematic analysis with the comparison with the Wasserstein-$2$ distance loss function, will be investigated in future work.
@@
Suppose $G$ is a linear transport map function such that
@@
We remark that the Cauchy distribution does not have the finite second moment. Thus, the Wasserstein-$2$ distance is not well defined, i.e., $\mathrm{W}_{2}(p(\cdot,\gamma_{1}),p(\cdot,\gamma_{2}))=+\infty$.
@@ 6. Discussion
This paper proposes a class of transport alpha divergences, one-parameter variation of the transport KL divergence and the transport Hessian distance. They are connected with Hessian metrics and 33–symmetric tensors of the negative Boltzmann-Shannon entropy in Wasserstein-$2$ space. We provide several analytical examples in one-dimensional probability densities, including generative models and Cauchy distributions.
It is worth mentioning that the quantile density functions (QDFs) have been applied in statistical learning problems [29]. The quantile density functions measure densities’ shape up to any constant shifting. The transport alpha divergence provides a class of functionals for measuring the difference from QDFs, i.e., derivatives of transport map functions. In future work, we shall study transport alpha divergences in high dimensional probability densities [10, 20]. This direction includes analysis, dualities, invariance properties, and optimization algorithms of transport mapping-related divergence functionals. In particular, systematic geometric calculations for Hessian structures in Wasserstein-$2$ space $(\mathcal{P}_{o}(\Omega),g_{\mathrm{H}},T_{\mathrm{H}})$ will be investigated; see related studies in [9, 20, 21, 26]. The systematic understanding of three symmetric tensors and alpha connections rely on both studies in optimal transport and information geometry; see related studies in [5, 6, 24].
Acknowledgements. W. Li’s work is supported by AFOSR YIP award No. FA9550-23-1-0087, NSF RTG: 2038080, NSF DMS: 2245097 and the McCausland Faculty Fellow in University of South Carolina.
@@
In this section, we provide some necessary proofs, which are used in this paper.
@@ 6.1. Derivatives in Wasserstein-$2$ space
We recall that the Levi-Civita connection induced geodesics in $(\mathcal{P}_{o}(\Omega),g_{\mathrm{W}})$ satisfies
$\left\{\begin{aligned} &\partial_{t}p(t,x)+\partial_{x}(p(t,x)\partial_{x}\Phi(t,x))=0\\ &\partial_{t}\Phi(t,x)+\frac{1}{2}|\partial_{x}\Phi(t,x)|^{2}=0,\end{aligned}\right$.
where $p(0,x)=p(x)$ and $\partial_{t}p(0,x)=\sigma(x)=-\partial_{x}(p(x)\Phi^{\prime}(x))$.
We prove the result by induction. When $n=1$, we have
@@
These geometric formulas are derived based on the Riemannian Levi-Civita connection in density manifold $(\mathcal{P}_{o}(\Omega),g_{\mathrm{W}})$. They formulate classical Gamma calculuses; see details in [16, 17, 30]. We leave the studies of high-order derivatives of negative entropy in $(\mathcal{P}_{o}(\Omega),g_{\mathrm{W}})$ in high dimensional sample spaces in future works.
@@ 6.2. Transport Hessian distance
Here the infimum is taken among all smooth density paths $p\colon[0,1]\times\Omega\rightarrow\mathbb{R}$, which connects both initial and terminal time probability density functions $q$, $p\in\mathcal{P}(\Omega)$. Interestingly, the variational problem (16) admits a closed-form expression.
@@
$\mathrm{Dist}_{\mathrm{H}}(p,q)^{2}=\int_{0}^{1}\left|\log\frac{Q_{p}^{\prime}(u)}{Q_{q}^{\prime}(u)}\right|^{2}du=2\mathrm{D}_{\mathrm{T},0}(p\|q).$
@@
where the infimum is taken among all smooth transport map functions $T\colon[0,1]\times\Omega\rightarrow\Omega$ with $T(0,x)=x$ and $T(1,x)=T(x)$. Thus,
@@
where the infimum is taken among all smooth paths $Q^{\prime}_{p_{t}}\colon[0,1]^{2}\rightarrow\mathbb{R}$ with fixed initial and terminal time conditions. By using the Euler-Lagrange equation for variable $Q^{\prime}_{p_{t}}$, we show that the geodesic equation in transport Hessian metric satisfies
@@ 6.3. Transport KL divergence
We also briefly present the derivation of transport KL divergence. Consider a functional $\mathcal{F}\colon\mathcal{P}\rightarrow\mathbb{R}$. We define a class of Bregman divergences in Wasserstein-$2$ space by
@@
where $p$, $q\in\mathcal{P}(\Omega)$, $\frac{\delta}{\delta q(x)}$ is the $L^{2}$ first variation w.r.t. $q(x)$, and $T_{\#}q=p$. See examples of $L^{2}$ first variation of functionals in [3, 10.4.2]. We refer $\mathrm{D}_{\mathrm{T},\mathcal{F}}$ as the transport Bregman divergence. If $\mathcal{F}$ is a second moment functional, i.e. $\mathcal{F}(p)=\int_{\Omega}|x|^{2}p(x)dx$, then $\mathrm{D}_{\mathrm{T},\mathcal{F}}$ forms the Wasserstein-2 distance. If $\mathcal{F}(p)=-\mathcal{H}(p)=\int_{\Omega}p(x)\log p(x)dx$, then
@@
where we use the fact that $q(x)\partial_{x}\log q(x)=\partial_{x}q(x)$ in the third equality, and apply the integration by parts in the last equality. From now on, we name $\mathrm{D}_{\mathrm{TKL}}$ the transport KL divergence. Again, by the change of variable (18), we note that $\mathrm{D}_{\mathrm{TKL}}$ can be formulated in terms of quantile density functions:
@@
From this formulation, we observe that $\mathrm{D}_{\mathrm{TKL}}$ is an Itakura–Saito type divergence in term of transport map functions or quantile density functions.
@@ 6.4. Validation of canonical divergences
We follow the idea of proof in [7], with using the change of variable (18) and the transport Hessian metric $g_{\mathrm{H}}$. We denote a monotone function $\hat{T}(x)=T^{-1}(x)$, such that $(T^{-1})_{\#}p=q$. Denote $\hat{T}\colon[0,1]\times\Omega\rightarrow\Omega$ satisfying Proposition 4 with $\alpha=-1$, such that
@@
$\begin{split}g_{\mathrm{H}}(\partial_{t}r,\partial_{t}r)=&\int_{\Omega}\left|\partial_{t}\partial_{x}\hat{T}(t,x)\frac{1}{\partial_{x}\hat{T}(t,x)}\right|^{2}{q(x)}dx.\end{split}$
@@
$\begin{split}-\frac{d^{2}}{dt^{2}}\mathcal{H}(r(t,\cdot))=&\int_{\Omega}\frac{d^{2}}{dt^{2}}\log\frac{p(x)}{\partial_{x}\hat{T}(t,x)}p(x)dx\\ =&\int_{\Omega}\left(\left|\partial_{t}\partial_{x}\hat{T}(t,x)\frac{1}{\partial_{x}\hat{T}(t,x)}\right|^{2}-\frac{\partial_{tt}\partial_{x}\hat{T}(t,x)}{\partial_{x}\hat{T}(t,x)}\right){q(x)}dx\\ =&g_{\mathrm{H}}(\partial_{t}r,\partial_{t}r).\end{split}$